# NeRF-MAE: Masked AutoEncoders for Self-Supervised 3D Representation Learning for Neural Radiance Fields

### v2

@@ 4 Experiments
Baselines: We compare 7 model variants across 3 different tasks to show the effectiveness of our method. 1) NeRF-RPN [hu2023nerf]: A state-of-the-art approach that predicts 3D boxes directly in NeRFs using 3D FPN [lin2017feature]. 2) Point-MAE [pang2022masked]: Extends 2D MAE [he2021masked] to 3D pointclouds and improves object classification and segmentation. 3) Point-M2AE [zhang2022point]: Extends Point-MAE to add multiscale features for irregular point cloud representation learning. 4) DepthContrast [zhang_depth_contrast]: Uses depth to learn 3D representations in a contrastive manner by comparing transformations of a 3D point cloud. 5) Contrastive Scene Context [hou2021exploring]: Leverages point correspondences along with spatial contexts to learn 3D scene level representation. 6) ImVoxelNet [rukhovich2022imvoxelnet]: A strong multi-view 3D detection baseline. 7) NeRF-MAE: Our 3D representation learning approach utilizing posed 2D images as input and operates directly on scene’s volumetric radiance and density grids obtained through NeRFs.
Datasets: We pre-train our 3D representation in a self-supervised manner on Front3D [fu20213d], HM3D [ramakrishnan2021hm3d] and Hypersim [roberts2021hypersim]. For cross-dataset transfer 3D object detection, we test on hold-out Scannet [dai2017scannet] scenes, where we use splits provided by [hu2023nerf]. For further details, see Sec. 3.3 and supplementary.
@@ 4.2 Comparisons with strong baselines for improved downstream performance
3D Object Detection Downstream Task: The results of our proposed method are summarized in Tables 2 and 4. We consistently outperform the state-of-the-art baseline methods on 3D object detection in NeRFs on both Front3D and ScanNet scenes. Among our variants, our method pre-trained on multiple data sources (Table 2, row 5) and with augmentations enabled during pretraining performs the best. Specifically, NeRF-MAE shows superior performance on the unseen test-set by achieving an AP50 of 63% and a Recall50 of 74.3%, hence demonstrating an absolute improvement of 21.5% in AP50 and 12.1% in Recall50 on the Front-3D dataset. Our method also achieves superior performance on ScanNet which is never seen during pretraining. We achieved an AP50 of 17% and a Recall50 of 39.5%, hence demonstrating an absolute performance improvement of 3% AP50 and 7.2% on Recall50 over the best-performing baseline.
@@ 6 Conclusion
Neural Fields have shown emergent properties going beyond reconstruction; hence making them useful for inferring semantics, geometry, and dynamics. In this work, we primarily study scaling NeRFs for their self-supervised pretraining, through the use of masked auto-encoders. In essence, we use a single standard Transformer model to learn useful 3D representations from NeRF which we show to be highly effective for 3D transfer learning. To pre-train our representation, we curate a large-scale dataset totaling 1.6M+ images and 3,500+ scenes from 4 different sources. Our approach, NeRF-MAE significantly outperforms self-supervised 3D pretraining as well as NeRF scene understanding baselines on a variety of downstream tasks. As it requires readily available and cheaper to obtain posed-RGB data and with a strong empirical quantitative performance over baselines, we hope this avenue is a potential future direction of 3D pretraining.
@@ C Experiment Analysis
In this section, we present additional experimental analysis. First, we present in detail an additional discussion on the downstream 3D object detection, semantic voxel-labeling, and voxel super-resolution tasks in Section C.1. We present the additional analysis of scaling laws and ablation studies in Section C.2 specifically highlighting the architectural ablation of adding skip connections to the decoder and discussing masking ratio ablation. We present additional insights with comparison to 3D pretraining baselines in Section C.3.

### v3

@@ 4 Experiments
Baselines: We compare 7 model variants across 3 different tasks to show the effectiveness of our method. 1) NeRF-RPN [hu2023nerf]: State-of-the-art method that predicts 3D boxes directly in NeRFs. 2) Point-MAE [pang2022masked]: Extends 2D MAE [he2021masked] to 3D pointclouds and improves classification and segmentation. 3) Point-M2AE [zhang2022point]: Extends Point-MAE to add multiscale features for point cloud representation learning. 4) DepthContrast [zhang_depth_contrast]: Uses depth to learn 3D representations in a contrastive manner by comparing transformations of a 3D point cloud. 5) Contrastive Scene Context [hou2021exploring]: Leverages point correspondences along with spatial contexts to learn 3D representations. 6) ImVoxelNet [rukhovich2022imvoxelnet]: A strong multi-view 3D detection baseline. 7) NeRF-MAE: Our 3D representation learning approach utilizing posed 2D images as input and operating directly on scene’s volumetric radiance and density grids obtained through NeRFs.
Datasets: We pretrain NeRF-MAE in a self-supervised manner on Front3D [fu20213d], HM3D [ramakrishnan2021hm3d] and Hypersim [roberts2021hypersim]. For cross-dataset transfer, we test on hold-out Scannet [dai2017scannet], on splits provided by [hu2023nerf]. For further details, see Sec. 3.3 and supplementary.
@@ 4.2 Comparisons with strong baselines for improved downstream performance
3D Object Detection Downstream Task: The results of our proposed method are summarized in Tables 2 and 4. We consistently outperform the state-of-the-art baseline methods on 3D object detection in NeRFs on both Front3D and ScanNet scenes. Among our variants, our method pretrained on multiple data sources (Table 2, row 5) and with augmentations enabled during pretraining performs the best. Specifically, NeRF-MAE shows superior performance on the unseen test-set by achieving an AP50 of 63% and a Recall50 of 74.3%, hence demonstrating an absolute improvement of 21.5% in AP50 and 12.1% in Recall50 on the Front-3D dataset. Our method also achieves superior performance on ScanNet which is never seen during pretraining. We achieved an AP50 of 17% and a Recall50 of 39.5%, hence demonstrating an absolute performance improvement of 3% AP50 and 7.2% on Recall50 over the best-performing baseline.
@@ 6 Conclusion
Neural Fields have shown emergent properties going beyond reconstruction; hence making them useful for inferring semantics, geometry, and dynamics. In this work, we use masked auto-encoders to study scaling NeRFs for their self-supervised pretraining. In essence, we use a single standard Transformer model to learn useful 3D representations from NeRF which we show to be highly effective for 3D transfer learning. To pretrain and evaluate our representation, we curate a large-scale dataset totaling 1.8M+ images and 3,600+ scenes from 4 different sources. Our approach, NeRF-MAE significantly outperforms self-supervised 3D pretraining as well as NeRF scene understanding baselines on a variety of downstream tasks. As it requires readily available and cheaper to obtain posed-RGB data and with a strong empirical quantitative performance over baselines, we hope this avenue is a potential future direction of 3D pertaining.
@@ B Implementation Details
Pretraining and Fine-tuning Weights: In the spirit of the original 2D MAE work [he2021masked], the pretraining weights of the model are general-purpose, containing the scene understanding knowledge such as semantic and spatial structure. Finetuning allows the model weights to be task-specific and aligns the model with the task using lightweight decoders. We will further include a detailed discussion on this topic in our paper.
@@ C Experiment Analysis
In this section, we present additional experimental analysis. First, we present in detail an additional discussion on the downstream 3D object detection, semantic voxel-labeling, and voxel super-resolution tasks in Section C.1. We present the additional analysis of scaling laws and ablation studies in Section C.2 specifically highlighting the architectural ablation of adding skip connections to the decoder and discussing masking ratio ablation. We present additional insights with comparison to 3D pretraining baselines in Section C.3 and details about per-scene NeRF optimization to prepare our pretraining dataset in Section C.4
@@ C.3 Comparison to 3D pretraining baselines
C.4 Per-scene NeRF optimization Analysis
Our per-scene NeRF optimization takes 8 mins on 1 GPU for a total of just over 2 days on 8 GPUs for the entire pre-training dataset. A by-product of this process that we would like to highlight is the compression of the original data (see Table 9). Our results indicate significant memory savings, and the resulting compressed representation still conveys all the necessary information for downstream tasks.
HM3D [ramakrishnan2021hm3d] (NeRF-MAE) ScanNet [dai2017scannet] test (NeRF-MAE)
[…]
Original (RGB) 256GB $\pm$ 0.00% 30GB $\pm$ 0.00%
Compressed (Weights) 46GB -- 82.03% 1.8GB -- 94.00%
Table 9: Relative Pretraining Dataset Comparison: showing significant memory saving during data preprocessing for NeRF-MAE.